There are four steps in the manufacturing process of a stuffed toy: cutting, stuffing, sealing, and packaging. There are two employees each for cutting and stuffing but one each for sealing and packaging. The processing times of cutting, stuffing, sealing, and packaging are 9, 4, 2, and 2.5 seconds per toy per employee.

a) What is the designed cycle time for stage "cutting" (in [min/unit])?

b) What is the designed cycle time for stage "sealing" (in [min/unit])?

c) Which stage is the bottleneck? What is the process capacity in [units/hour]? If you add one employee in parallel at the bottleneck stage, what should be the capacity of the new employee so that you increase the process capacity as much as possible?

thank you.
Answer #1

a)

Ccle time is the average time between completing the task. For cutting we have 2 employees and each can complete the task in 9 seconds. This mean 1 min we can make 2*60/9 = 13.33 units. This means the cycle time is 1/13.33 = 0.075 min/unit

b)

In 1 min, we can seal 60/2 = 30 units. This means the cycle time 1/30 = 0.033 min/unit

c)

The capacity at each stage in units/hour

Cutting = 2*(60*60/9) = 800 units

Stuffing = 2*(60*60)/4 = 1800 units

Sealing = (60*60)/2 = 1800 units

Packaging = (60*60)/2.5 = 1440 units

The system capacity = min(800,1800,1800,1440) = 800 unit/hour

The bottleneck stage is cutting. If we add one more employee at the cutting stage the capacity will be 3*(60*60)/9 = 1200 units per hour